OpenStudy (anonymous):

how to solve (x/2)-1=(x/5)+2

OpenStudy (anonymous):

x/2 - x/5 = 3 change denominator to 10: 5x/10 - 2x/10 = 3 3x/10 = 3 3x = 30 x = 10